What concentration of NH3 is required to make a solution with a pH of 11.22.



1
Expert's answer
2021-03-25T03:41:50-0400

Solution:

Kb for NH3 = 1.8×10-5

From the pH: pOH = 14 - pH = 14 - 11.22 = 2.78

Hence,

[OH-] = 10-pOH = 10-2.78 = 0.00166

[OH-] = 0.00166 M


NH3 + H2O = NH4+ + OH-

Hence,

[NH4+] = [OH-] = 0.00166 M.

[NH3] = (x - 0.00166) M


Kb = [NH4+][OH-]/[NH3]

Hence,

1.8×10-5 = (0.00166)2 / (x - 0.00166)

Assume that x >> 0.00166

1.8×10-5 = (0.00166)2 / (x)

1.8×10-5x = 2.756×10-6

x = [NH3] = 0.153 M

0.153 is >> 0.00166


Answer: [NH3] = 0.153 M